A container of top soil was found to weigh 40.0 kg and measures 14.0cm×20.0cm×40.0cm. What is the specific geavity of top soil?
1
Expert's answer
2020-06-26T01:03:28-0400

Density of top soil = (40.0×1000) /(14.0×20.0×40.0) g/cm3

                            = 3.571 g/cm3

Specific gravity = density of the object / density of pure water   [density of pure water =1 g/cm3]

                         = 3.571/1 =3.571 (Answer)